Compute the relative rate of diffusion of ethanol (C2H5OH) to propane (C3H8)
1
Expert's answer
2014-12-23T10:58:26-0500
The relative rate of diffusion will be sqrt (M(C3H8)/M(C2H5OH)) = sqrt (44/46) = 0.978